# Let Me Teach You: Pedagogical Foundations of Feedback for Language Models

### v2

@@ 1 Introduction
The last few years introduced a new paradigm for finetuning Large Language Models (LLMs): learning from human feedback (Ziegler et al., 2020; Stiennon et al., 2022; Bai et al., 2022a; OpenAI, 2023; Rafailov et al., 2023; Azar et al., 2023; Fisch et al., 2024) to augment their capabilities beyond their pretraining (Christiano et al., 2017; Wu et al., 2021; Menick et al., 2022). This alignment has yielded less toxic and harmful models (Bai et al., 2022b; Korbak et al., 2023) that are preferred by users (Ouyang et al., 2022a; Bai et al., 2022b).
@@
Whether learning from feedback is done by directly learning from human preferences (Rafailov et al., 2023; Azar et al., 2023; Hong et al., 2024; Meng et al., 2024; Saeidi et al., 2024; Fisch et al., 2024), or with Reinforcement Learning from Feedback (RLF) using human-curated (RLHF; Ouyang et al., 2022a; Bai et al., 2022a; Ramamurthy et al., 2023) or AI-generated (RLAIF; Bai et al., 2022b; Saunders et al., 2022; Madaan et al., 2023) feedback, all variants have been shown to be successful in several metrics — from encouraging honest behaviors, to reducing toxicity, to being preferred by evaluators (Ouyang et al., 2022a). Other approaches, such as imitation learning (Li et al., 2016; Kreutzer et al., 2018; Hancock et al., 2019; Scheurer et al., 2022), and feedback modeling (Weston, 2016; Li et al., 2017; Hancock et al., 2019; Xu et al., 2022; Liu et al., 2023) have had similar success. Feedback has thus emerged as an important source of information for model improvement and evaluation, directing them toward desired objectives and behaviors (Fernandes et al., 2023).
@@ 2.1 Current State of Feedback in NLP
The value of feedback is derived from the implicit knowledge it represents about human values and expectations, that would otherwise be extremely difficult to specify (Christiano et al., 2017). Feedback can assume different forms: numerical ratings, rankings, preferences, demonstrations, and textual information (either using a template or free-form text — structured and unstructured feedback, respectively). RLF methodologies usually collect either a numeric rating or a ranking for classifying the quality of an initial answer (often encouraging properties such as helpfulness and honesty while mitigating harmfulness; Askell et al., 2021). RLF may also leverage demonstrations to finetune LLMs in a supervised fashion before the RLF stage to reduce the search space (Ouyang et al., 2022a; Bai et al., 2022b). RLF methods address the intractable problem of designing an appropriate loss function for training models to exhibit behaviors with no closed-form solution (Askell et al., 2021). Recent works have also started to incorporate feedback in a prompt-based manner (Schick et al., 2022; Madaan et al., 2023; Paul et al., 2023; Chen et al., 2023; Lin et al., 2023; Zhao et al., 2024).
@@
However, the extent to which feedback formulations transmit their goal states remains unclear. For example, InstructGPT (Ouyang et al., 2022b) is first finetuned on demonstration data, and then performs RLHF with a reward model trained using comparison data (i.e., pairs of ranked generations). This feedback is limited in the information it transmits. Scoring a demonstration A as “better” than demonstration B provides little information on the quality of A nor B,  [^1] We note such a format also obfuscates any bias and disagreement that occurred in reaching such a judgment. nor on how either demonstration can be improved. Taking these limitations into account, RMs are likely to suffer from some degree of misalignment (Pan et al., 2022; Gao et al., 2022; Song et al., 2023). Recent works have started to acknowledge the limited information in these feedback formulations, recognizing them as unsuited for capturing critical information, such as different types of errors (Golovneva et al., 2023; Wu et al., 2023).
@@ 2.2 Natural Language Feedback for LLMs
Certain works augment the model through data augmentation (Shi et al., 2022), external corrective feedback (Tandon et al., 2022; Madaan et al., 2022; Shinn et al., 2023) or natural language patches (Murty et al., 2022). Another line of work introduces a secondary model, that either refines an original LLM’s answer (Scheurer et al., 2022; Welleck et al., 2022; Tandon et al., 2022), critiques it (Saunders et al., 2022; Paul et al., 2023) iteratively self-improves (Schick et al., 2022; Chen et al., 2023; Madaan et al., 2023; Yuan et al., 2024) or otherwise constrains the initial response of a model (Stephan et al., 2024). Beyond increasing the complexity of feedback using natural language, several of these approaches also target intermediate generations with their feedback, not the final outcome produced by the model, thereby increasing the number of feedback opportunities through multiple iterations (Lightman et al., 2023).
@@ 3.2 What constitutes effective feedback?
Kluger and DeNisi (1996) showed feedback was detrimental to a learner’s performance in $38\%$ of analyzed cases. Three main requirements for helpful feedback emerge from previous work: applicability, learner regulation, and personalization. These requirements are directly related to feedback content, which we explore in §4.2 for LLMs.
@@
Feedback should be actionable, allowing the learner to achieve a desired target performance. Sadler (1989) suggest that feedback needs to identify a target performance, compare the learner’s current performance to it, and engage in actions to reduce that difference. Similarly, Hattie and Timperley (2007) indicate that effective feedback needs to answer three questions: where the learner is going (the goal), how they can get there, and where to go next. Other works extend these definitions of effective feedback by including elements such as motivational and metacognitive aspects Nicol and Macfarlane-Dick (2006) or aspects of teaching (e.g., lesson design; Evans, 2013).
@@ 3.3 What are the characteristics of feedback?
Works in this category focus on the characteristics of the content only. Kulhavy and Stock (1989), for example, model feedback through a verification component, which is a simple discrete classification of the answer as correct or incorrect, and an elaboration component, which contains all other information. Other works Hattie and Timperley (2007); Panadero and Lipnevich (2022) suggest three categories for classifying feedback: (i) addressing the learner’s performance goal, (ii) addressing the learner’s current performance, and (iii) addressing the next steps the learner should undertake. Our feedback content taxonomy draws primarily from the analysis and adaptation these works to the NLP domain. Additionally, several categorizations include feedback about mistakes, from which we derive the Error component of FELT4).
@@ 3.3.2 What variables influence feedback?
Mason and Bruning (2001) consider student achievement and prior knowledge as important factors impacting feedback optimality. Nicol and Macfarlane-Dick (2006) expands prior knowledge and proficiency into domain knowledge and strategy knowledge (along with motivational beliefs), updated upon each attempt through both internal and external feedback. Narciss and Huth (2004); Narciss (2008) flesh out the learner characteristics further, including learning goals and motivation. Lipnevich et al. (2016) identify “personality, general cognitive ability, receptivity to feedback, prior knowledge, and motivation” as key learner characteristics that impact feedback processing.
Other works focus on the learner’s feedback processing mechanisms. Kluger and DeNisi (1996) propose three processing levels: details about how to solve the task, the task itself as a whole, and meta-task processes. In contrast, Hattie and Timperley (2007) present four levels of feedback processing: (1) task level, how well the tasks are understood and performed, (2) process level, the process needed to achieve the task level understanding and performance, (3) self-regulation level, the self-monitoring and the direction and regulation of the learner’s actions, and (4) self level, personal evaluations about the learner as a whole.
@@ 4 FELT: Unifying The Two Worlds
In the pedogogical research surveyed in the previous section (§3), feedback emerges as both a complex ecosystem, and a rich, but systematized, information source. In this section, we consolidate different pedagogical feedback models, and conceptualize a novel feedback framework adapted for LLMs, FELT (Feedback, Errors, Learner, Task), displayed in Figure 2. We briefly analyze each of its components, and dedicate a specific section (§4.2) to the formalization of feedback content.
@@ 4.1 The Big Picture
The learner in FELT represents the LLM itself, represented by two components: prior knowledge (which is dependent on the task), and the feedback processing mechanism. Prior knowledge is captured by the model’s size, pretraining data and pretraining method. Model size is suggested to be linked to the model’s ability to effectively leverage feedback (Scheurer et al., 2022; Bai et al., 2022b). The pretraining data, as well as how the model was trained, similarly encode its initial ability to tackle a given task. The feedback processing mechanism reflects the method used to transmit feedback to the model, naturally influencing how the feedback is subsequently processed. For example, training objectives necessarily influence how models process and incorporate feedback. Fernandes et al. (2023) identify three common feedback integration mechanisms: feedback-based imitation learning, joint-feedback modeling, and reinforcement learning (which can be generally extended to include other recent training methods, such as DPO, Rafailov et al., 2023). In addition, we also consider feedback provided using in-context learning (ICL; Brown et al., 2020; Madaan et al., 2022; Zhao et al., 2024). Much like the Task, the Learner (i.e., the models and algorithms used for learning from feedback) is an active area of NLP research.
@@
Feedback may relay information on where the learner is failing, which requires understanding the possible error modes for a given task, and which ones the learner is likely in. For example, guessing and committing systematic reasoning mistakes reflect differing degrees of understanding. Moreover, not all errors should be treated the same. Wu et al. (2023) find that assigning distinct reward models to specific error modes (e.g., conciseness, factuality, relevance) improved performance over using a single RM, clarifying the importance of modeling the error space – another active area of research (Golovneva et al., 2023; Paul et al., 2023; Murugesan et al., 2023; Mishra et al., 2024).
@@
Timing specifies whether feedback is provided immediately or after a delay. Certain types of prompting can be formulated as variations in timing. For example, Metacognitive reasoning (Wang and Zhao, 2023) prompts the model to answer a question, critically evaluate its answer, possibly revise it, and assess its confidence in it. In contrast, Chain-of-Thought (CoT; Wei et al., 2022) does not elicit any post-answer reasoning. Coupled with feedback learning methods such as RLF or DPO, this “padding” between the answer and the reward signal might affect which prompting strategy leads to the best performance. Delayed feedback also has the potential for deriving multiple model updates from a single initial data point. From an answer and a reflection generated by an LLM (i.e., an attempt to identify and correct errors on that answer), and a piece of feedback, a model can learn not only whether the answer is correct, but also whether its reflection is adequate — applicable both in ICL and traditional training (e.g., RLF, DPO) settings.
@@ 5 Realizing the Promise of Feedback
FELT provides a springboard for exploring the entire feedback ecosystem. To understand what makes feedback effective for LLMs, we briefly propose some ideas for future investigation (and categorize ideas from previous studies) that demonstrate the important role that each of the different components of FELT play in this exploration.
Task. Feedback might influence model performance differently depending on the task formulation for the LLM (e.g., asking the model to extract entities from text as opposed to asking the model to simply identify their start and end positions).
Learner. To explore the degree to which alignment extends beyond changing the linguistic style of AI models (Lin et al., 2023), the pretraining and finetuning data of LLMs can be modified to assess the impact feedback has on a model both with and without prior knowledge. Uncertainty awareness has also been show to improve model alignment (Wang et al., 2024). Exploring the model’s receptiveness to feedback conditioned on its uncertainty may enable better understanding of what leads models to reject their own parametric knowledge.
Errors. Implementing different penalties for various error types will lead to different model behaviors (Wu et al., 2023). This flexibility will yield a better understanding of which degrees of penalization lead to a given behavior, as well as how much this penalization should be aligned with the actual severity that the error merits.
@@
A limitation of prior works on NLF is the indiscriminate treatment of different types of feedback. [^9] Even in the same work, authors might treat feedback such as “write a step-by-step reasoning before the solution” as equivalent to “should be contextually relevant and easy to read.” As made evident by our taxonomy, they are, however, clearly different, as the former provides procedural instructions and the latter peripheral information clarifying the task. The four types of feedback content — learner status, goal, procedural, and peripheral — directly address this issue, allowing more systematic studies of feedback.
@@
ICL is an active area of research in NLP, including in model alignment (Lin et al., 2023; Zhao et al., 2024). Several dimensions of our feedback content taxonomy have been shown to contribute to either aligning LLMs to specific desired behaviors, or to increasing its performance in a given task — suggesting both the utility of these content dimensions, and hinting at many as of yet unexplored approaches that vary them in novel ways. Prompts with a high degree of applicability of instructions and with a focus on procedural types of information are common, be it asking for a rationale (Wei et al., 2022; Wang and Zhao, 2023; Yao et al., 2023), or more task-specific instructions (Wang and Zhao, 2023; Madaan et al., 2023). Prompting the model with goal information is also popular, most frequently done using few-shot prompting (Brown et al., 2020). Valence is often introduced either from (explicit or implicit) human feedback (Shi et al., 2022; Pang et al., 2024) or by criticizing a preliminary answer (Bai et al., 2022b; Paul et al., 2023; Yao et al., 2023; Yuan et al., 2024), providing the model with information on its learner status. The style, or linguistic properties, of prompts can lead to significant performance variations (Leidinger et al., 2023). Finally, while Information Novelty is hard to measure for LLMs, recent research has investigated what causes LLMs to be faithful to the novel information rather than their encoded knowledge (Longpre et al., 2021), an important quality to understand and predict model behavior (Zhou et al., 2023; Yin et al., 2023). All these dimensions underlie a wide range of instruction formulations for LLMs, enabling practitioners to taxonomize existing prompting strategies, and uncover novel approaches based on different combinations.

### v3

@@ 1 Introduction
The last few years introduced a new paradigm for finetuning Large Language Models (LLMs): learning from human feedback (Ziegler et al., 2020; Stiennon et al., 2020; Bai et al., 2022a; OpenAI, 2023; Rafailov et al., 2023; Gheshlaghi Azar et al., 2024; Fisch et al., 2024) to augment their capabilities beyond their pretraining (Christiano et al., 2017; Wu et al., 2021; Menick et al., 2022). This alignment has yielded less toxic and harmful models (Bai et al., 2022b; Korbak et al., 2023) that are preferred by users (Ouyang et al., 2024; Bai et al., 2022b).
@@
Whether learning from feedback is done by directly learning from human preferences (Rafailov et al., 2023; Gheshlaghi Azar et al., 2024; Hong et al., 2024; Meng et al., 2024; Saeidi et al., 2024; Fisch et al., 2024), or with Reinforcement Learning from Feedback (RLF) using human-curated (RLHF; Ouyang et al., 2024; Bai et al., 2022a; Ramamurthy et al., 2023) or AI-generated (RLAIF; Bai et al., 2022b; Saunders et al., 2022; Madaan et al., 2023) feedback, all variants have been shown to be successful in several metrics — from encouraging honest behaviors, to reducing toxicity, to being preferred by evaluators (Ouyang et al., 2024). Other approaches, such as imitation learning (Li et al., 2017; Kreutzer et al., 2018; Hancock et al., 2019; Scheurer et al., 2022), and feedback modeling (Weston, 2016; Li et al., 2017; Hancock et al., 2019; Xu et al., 2023; Liu et al., 2023) have had similar success. Feedback has thus emerged as an important source of information for model improvement and evaluation, directing them toward desired objectives and behaviors (Fernandes et al., 2023).
@@ 2.1 Current State of Feedback in NLP
The value of feedback is derived from the implicit knowledge it represents about human values and expectations, that would otherwise be extremely difficult to specify (Christiano et al., 2017). Feedback can assume different forms: numerical ratings, rankings, preferences, demonstrations, and textual information (either using a template or free-form text — structured and unstructured feedback, respectively). RLF methodologies usually collect either a numeric rating or a ranking for classifying the quality of an initial answer (often encouraging properties such as helpfulness and honesty while mitigating harmfulness; Askell et al., 2021). RLF may also leverage demonstrations to finetune LLMs in a supervised fashion before the RLF stage to reduce the search space (Ouyang et al., 2024; Bai et al., 2022b). RLF methods address the intractable problem of designing an appropriate loss function for training models to exhibit behaviors with no closed-form solution (Askell et al., 2021). Recent works have also started to incorporate feedback in a prompt-based manner (Schick et al., 2023; Madaan et al., 2023; Paul et al., 2024; Chen et al., 2024; Lin et al., 2024; Zhao et al., 2024).
@@
However, the extent to which feedback formulations transmit their goal states remains unclear. For example, InstructGPT (Ouyang et al., 2024) is first finetuned on demonstration data, and then performs RLHF with a reward model trained using comparison data (i.e., pairs of ranked generations). This feedback is limited in the information it transmits. Scoring a demonstration A as “better” than demonstration B provides little information on the quality of A nor B,  [^1] We note such a format also obfuscates any bias and disagreement that occurred in reaching such a judgment. nor on how either demonstration can be improved. Taking these limitations into account, RMs are likely to suffer from some degree of misalignment (Pan et al., 2022; Gao et al., 2023; Song et al., 2023). Recent works have started to acknowledge the limited information in these feedback formulations, recognizing them as unsuited for capturing critical information, such as different types of errors (Golovneva et al., 2023; Wu et al., 2023).
@@ 2.2 Natural Language Feedback for LLMs
Certain works augment the model through data augmentation (Shi et al., 2024), external corrective feedback (Tandon et al., 2022; Madaan et al., 2022; Shinn et al., 2023) or natural language patches (Murty et al., 2022). Another line of work introduces a secondary model, that either refines an original LLM’s answer (Scheurer et al., 2022; Welleck et al., 2023; Tandon et al., 2022), critiques it (Saunders et al., 2022; Paul et al., 2024) iteratively self-improves (Schick et al., 2023; Chen et al., 2024; Madaan et al., 2023; Yuan et al., 2024) or otherwise constrains the initial response of a model (Stephan et al., 2024). Beyond increasing the complexity of feedback using natural language, several of these approaches also target intermediate generations with their feedback, not the final outcome produced by the model, thereby increasing the number of feedback opportunities through multiple iterations (Lightman et al., 2024).
@@ 3.2 What constitutes effective feedback?
Kluger and DeNisi (1996) showed feedback was detrimental to a learner’s performance in $38\%$ of analyzed cases. Three main requirements for helpful feedback emerge from previous work: applicability, learner regulation, and personalization. These requirements are directly related to feedback content, which we explore in §4.2 for LLMs. [^3] Looking forth, Applicability corresponds to dimensions of applicability of instructions, learner regulation to purpose, and personalization is a product of the combination of all dimensions, allowing a fine-grained customization of feedback.
@@
Feedback should be actionable, allowing the learner to achieve a desired target performance. Sadler (1989) suggests that feedback needs to identify a target performance, compare the learner’s current performance to it, and engage in actions to reduce that difference. Similarly, Hattie and Timperley (2007) indicate that effective feedback needs to answer three questions: where the learner is going (the goal), how they can get there, and where to go next. Other works extend these definitions of effective feedback by including elements such as motivational and metacognitive aspects Nicol and Macfarlane-Dick (2006) or aspects of teaching (e.g., lesson design; Evans, 2013).
@@ 3.3 What are the characteristics of feedback?
Works in this category focus on the characteristics of the content only. Kulhavy and Stock (1989), for example, model feedback through a verification component, which is a simple discrete classification of the answer as correct or incorrect, and an elaboration component, which contains all other information. Other works Hattie and Timperley (2007); Panadero and Lipnevich (2022) suggest three categories for classifying feedback: (i) addressing the learner’s performance goal, (ii) addressing the learner’s current performance, and (iii) addressing the next steps the learner should undertake. Our feedback content taxonomy draws primarily from the analysis and adaptation of these works to the NLP domain. Additionally, several categorizations include feedback about mistakes, from which we derive the Error component of FELT4).
@@ 3.3.2 What variables influence feedback?
Mason and Bruning (2001) consider student achievement and prior knowledge as important factors impacting feedback optimality. Nicol and Macfarlane-Dick (2006) expand prior knowledge and proficiency into domain knowledge and strategy knowledge (along with motivational beliefs), updated upon each attempt through both internal and external feedback. Narciss and Huth (2004); Narciss (2008) flesh out the learner characteristics further, including learning goals and motivation. Lipnevich et al. (2016) identify “personality, general cognitive ability, receptivity to feedback, prior knowledge, and motivation” as key learner characteristics that impact feedback processing.
Other works focus on the learner’s feedback processing mechanisms. Kluger and DeNisi (1996) propose three processing levels: details about how to solve the task, the task itself as a whole, and meta-task processes. In contrast, Hattie and Timperley (2007) present four levels of feedback processing: (1) task level, how well the tasks are understood and performed, (2) process level, the process needed to achieve the task level understanding and performance, (3) self-regulation level, the self-monitoring and the direction and regulation of the learner’s actions, and (4) self level, personal evaluations about the learner as a whole (which are usually uninformative about the task and can be very hard to change, e.g., the quality of being “a good student”). Current LLMs do not seem capable of more metacognitive levels of feedback processing.
@@ 4 FELT: Unifying The Two Worlds
In the pedagogical research surveyed in the previous section (§3), feedback emerges as both a complex ecosystem, and a rich, but systematized, information source. In this section, we consolidate different pedagogical feedback models, and conceptualize a novel feedback framework adapted for LLMs, FELT (Feedback, Errors, Learner, Task), displayed in Figure 2. We briefly analyze each of its components, and dedicate a specific section (§4.2) to the formalization of feedback content.
@@ 4.1 The Big Picture
The learner in FELT represents the LLM itself, represented by two components: prior knowledge (which is dependent on the task), and the feedback processing mechanism. [^7] Several pedagogical dimensions for human feedback processing, discussed in §3, are not currently applicable to LLMs — such as the learner’s regulation, the learner’s affective processes and multiple motivations (e.g., both succeeding at the task and being seen as competent, which can be a deterrent to success if the learner believes they might fail), the learner’s perception of their control over the task, autonomous receptivity to feedback, and higher levels of feedback processing, including meta-task and self level learning — and as such do not have a direct representation in FELT’s current form. Prior knowledge is captured by the model’s size, pretraining data and pretraining method. Model size is suggested to be linked to the model’s ability to effectively leverage feedback (Scheurer et al., 2022; Bai et al., 2022b). The pretraining data, as well as how the model was trained, similarly encode its initial ability to tackle a given task. The feedback processing mechanism reflects the method used to transmit feedback to the model, naturally influencing how the feedback is subsequently processed. For example, training objectives necessarily influence how models process and incorporate feedback. Fernandes et al. (2023) identify three common feedback integration mechanisms: feedback-based imitation learning, joint-feedback modeling, and reinforcement learning (which can be generally extended to include other recent training methods, such as DPO, Rafailov et al., 2023). In addition, we also consider feedback provided using in-context learning (ICL; Brown et al., 2020; Madaan et al., 2022; Zhao et al., 2024). Much like the Task, the Learner (i.e., the models and algorithms used for learning from feedback) is an active area of NLP research.
@@
Feedback may relay information on where the learner is failing, which requires understanding the possible error modes for a given task, and which ones the learner is likely in. For example, guessing and committing systematic reasoning mistakes reflect differing degrees of understanding. Moreover, not all errors should be treated the same. Wu et al. (2023) find that assigning distinct reward models to specific error modes (e.g., conciseness, factuality, relevance) improved performance over using a single RM, clarifying the importance of modeling the error space – another active area of research (Golovneva et al., 2023; Paul et al., 2024; Murugesan et al., 2023; Mishra et al., 2024).
@@
Timing specifies whether feedback is provided immediately or after a delay. Certain types of prompting can be formulated as variations in timing. For example, Metacognitive reasoning (Wang and Zhao, 2024) prompts the model to answer a question, critically evaluate its answer, possibly revise it, and assess its confidence in it. In contrast, Chain-of-Thought (CoT; Wei et al., 2024) does not elicit any post-answer reasoning. Coupled with feedback learning methods such as RLF or DPO, this “padding” between the answer and the reward signal might affect which prompting strategy leads to the best performance. Delayed feedback also has the potential for deriving multiple model updates from a single initial data point. From an answer and a reflection generated by an LLM (i.e., an attempt to identify and correct errors on that answer), and a piece of feedback, a model can learn not only whether the answer is correct, but also whether its reflection is adequate — applicable both in ICL and traditional training (e.g., RLF, DPO) settings.
@@ 5 Realizing the Promise of Feedback
Figure 3: Visual summary of several future research directions motivated by FELT, for each of its dimensions and across several research axes. The textual counterparts of each bubble are highlighted in their respective color (green, blue, purple, orange) in the main body of Section 5.
FELT provides a springboard for exploring the entire feedback ecosystem. To understand what makes feedback effective for LLMs, we briefly propose some ideas for future investigation (and categorize ideas from previous studies) that demonstrate the important role that each of the different components of FELT play in this exploration. Figure 3 presents a visual summary of the proposals presented in this section.
Task. Feedback might influence model performance differently depending on the goal formulation for the LLM (e.g., asking the model to extract entities from text v.s. asking the model to identify their start and end positions). Task-specific processing (e.g., LoRA; Hu et al., 2022) would allow models to adapt to a wider range of tasks.
Learner. To explore the degree to which alignment extends beyond changing the linguistic style of AI models (Lin et al., 2024), LLMs can be modeled as different types of learners — whether by (1) controlling their knowledge capacity, i.e., creating learners both with and without prior knowledge through data interventions during pretraining or finetuning, or external augmentation; and (2) implementing novel update mechanisms that can capture richer feedback (e.g., TextGrad; Yuksekgonul et al., 2024). Uncertainty awareness has also been show to improve model alignment (Wang et al., 2024). Exploring the model’s receptiveness to feedback conditioned on its uncertainty may enable better understanding of what leads models to reject their own parametric knowledge.
Errors. After mapping the different error types a model might make, it becomes possible to implement different penalties for these error types, which might promote different model behaviors (Wu et al., 2023). This flexibility will yield a better understanding of which degrees of penalization lead to a given behavior, as well as the degree to which this penalization should be aligned with the actual severity that the error merits.
@@
A limitation of prior works on NLF is the indiscriminate treatment of different types of feedback (shown in Table 1). The four types of feedback content — learner status, goal, procedural, and peripheral — directly address this issue, allowing more systematic studies of feedback. The natural next step is to provide models with various kinds of feedback and compare the impact of the different types of information and of its delivery.
Content Type Previous Works
Learner Status Weston (2016), Tandon et al. (2022), Scheurer et al. (2022), Saunders et al. (2022), Shi et al. (2024), Welleck et al. (2023), Wu et al. (2023), Shinn et al. (2023), Chen et al. (2024), Paul et al. (2024), Madaan et al. (2023), Lightman et al. (2024), Yoon et al. (2024)
Goal Weston (2016), Saunders et al. (2022), Shi et al. (2024)
Procedural Weston (2016), Tandon et al. (2022), Saunders et al. (2022), Shi et al. (2024), Welleck et al. (2023), Madaan et al. (2023), Murty et al. (2022), Schick et al. (2023)
Peripheral Tandon et al. (2022), Scheurer et al. (2022), Chen et al. (2024), Madaan et al. (2023)
Table 1: Past works using textual or otherwise augmented feedback. Even in the same work, feedback such as “write a step-by-step reasoning trace before the solution” is often treated as equivalent to “should be contextually relevant and easy to check.” As made evident by our taxonomy, however, they are markedly different, as the former provides procedural instructions and the latter peripheral information clarifying the task.
@@
ICL is an active area of research in NLP, including in model alignment (Lin et al., 2024; Zhao et al., 2024). Several dimensions of our feedback content taxonomy have been shown to contribute to either aligning LLMs to specific desired behaviors, or to increasing its performance in a given task — suggesting both the utility of these content dimensions, and hinting at many as of yet unexplored approaches that vary them in novel ways. Prompts with a high degree of applicability of instructions and with a focus on procedural types of information are common, be it asking for a rationale (Wei et al., 2024; Wang and Zhao, 2024; Yao et al., 2023), or more task-specific instructions (Wang and Zhao, 2024; Madaan et al., 2023). Prompting the model with goal information is also popular, most frequently done using few-shot prompting (Brown et al., 2020).
Valence is often introduced either from (explicit or implicit) human feedback (Shi et al., 2024; Pang et al., 2024) or by criticizing a preliminary answer (Bai et al., 2022b; Paul et al., 2024; Yao et al., 2023; Yuan et al., 2024), providing the model with information on its learner status. The style, or linguistic properties, of prompts can lead to significant performance variations (Leidinger et al., 2023). Finally, while Information Novelty is hard to measure for LLMs, recent research has investigated what causes LLMs to be faithful to the novel information rather than their encoded knowledge (Longpre et al., 2021), an important quality to understand and predict model behavior (Zhou et al., 2023; Yin et al., 2023). All these dimensions underlie a wide range of instruction formulations for LLMs, enabling practitioners to taxonomize existing prompting strategies, and uncover novel approaches based on different combinations.